n a jump spike, Mark Angel slams the volleyball from a height of 2.3 m with a speed of 18 m/s at 24° below the horizontal. How far on the opposite side will the ball land?


1
Expert's answer
2021-12-07T09:53:54-0500
"d=\\\frac{18\cos{24}}{9.8}(18\sin{24}+\sqrt{(18\sin{24})^2+2(9.8(2.3)})\\=29\ m"